OpenStudy (anonymous):

If cot x = .78, what is tan x?

OpenStudy (anonymous):

tanx=1/cot x

OpenStudy (amistre64):

what relation do cot and tan have? ...

OpenStudy (anonymous):

i have no idea what you guys are talking about ): im soooo bad at math/:

OpenStudy (anonymous):

tanx=1/cot x=1/.78

OpenStudy (anonymous):

they are saying if you have the value of cot them divide 1 by the value of cot to get tangent.cot is the reciprocal of tangent \[\Large \tan(\theta)=\frac{1}{\cot(\theta)}\]